# JWST Reveals Varied Origins Between Jupiter’s Irregular Satellites

### v1

@@ 7 Trojan-like Surfaces Exist Amongst Fragmented Irregular Satellites
Previous compositional assessment of the spectrally “red” irregular satellites have primarily relied on analogies to Jovian Trojan asteroids without specific mineralogical constraints. VNIR measurements (Sharkey et al., 2023; Grav & Holman, 2004) showed that several irregular satellites closely match Jovian Trojan color taxa. Figure 9 demonstrates color and absorption band similarities between the irregular satellites Sinope and Carme and the Jovian Trojans Leucus, Patroclus, and Eurybates (Wong et al., 2024). Sinope and Leucus are of comparable sizes (D 40 km) and have very similar spectra. Carme is an order of magnitude smaller than Patroclus. Carme and Patroclus have similar colors longward of 1.6 , but VNIR data discriminates Carme into the “red” taxon and differentiates it from the “less-red” Patroclus (Sharkey et al., 2019). This similarity in bands suggests that less-red materials could be present on redder surfaces.
The red objects (Carme, Sinope, Themisto, Figure 9, left) display similar 3.4 organics features with band depths of , consistent with Wong et al. (2024)’s suggested correlation between color and organic content. The red objects (Carme, Sinope, Themisto) have similar organic band depths of , while Patroclus’s band depth is . Carme may be consistent with less-red material that includes higher reddening from organics. Carme’s band properties suggest that it could be a transition object between red and less-red groups, or that it contains a mixture of both materials.
@@
The presence of Trojan-like 3.0 and 3.4 absorption bands on small red Jovian irregular satellites places some new constraints on the features’ origins. The irregular satellites’ extensive collisional processing (Bottke et al., 2010) took place after their capture about Jupiter. This post-capture processing suggests that irregular satellites could have younger surfaces compared to similarly sized Jovian Trojans (as Trojans larger than are unlikely to have been substantially collisionally evolved since their own capture, Nesvorný 2018; Marschall et al. 2022; Bottke et al. 2023). In order for the same 3.0 and 3.4 absorbers to be present across a variety of irregular satellite sizes (), these materials must have been abundant enough in irregular satellite parent bodies to produce red collisional fragments (for example, if red objects have red interiors). Alternatively, red materials could be produced on satellite surfaces faster than their surfaces are collisionally refreshed. If red surfaces are formed in-place, then their formation must be related to materials that are stable within the Jovian thermal environment over those weathering timescales.
The less-red objects Pasiphae and Lysithea have 3 features which are distinct and shifted to shorter wavelengths (near ) compared to the redder objects Leucus and Sinope (bands near ). Fig. 10 shows that the band shapes observed on Lysithea and Pasiphae are good matches to those observed on Eurybates and Patroclus (Wong et al., 2024). The 3 band depths of Patroclus, Eurybates, and Lysithea/Pasiphae do not display a trend as a function of their VNIR colors.
The similarities between the absorptions on Eurybates and Pasiphae/Lysithea is particularly striking given Eurybates’ uniqueness amongst Trojans. The Eurybates collisional family is more spectrally neutral than typical Trojans and is similar in spectral slope to C-type asteroids (Roig et al., 2008). As a major outlier, it has been suggested that Eurybates may demonstrate what a younger, non-weathered Trojan surface may look like. The irregular satellites demonstrate that satellites of the same size in the same family can express a wide range of Trojan-like spectra. Figure 10 also compares the spectra of Lysithea and Eurybates to the dark, ice-poor regions of Phoebe as observed by the Cassini mission (Clark et al., 2005). This connection suggests that captured bodies at Jupiter and Saturn have similar non-ice components.
The origins of the absorber(s) among the less-red objects are also unclear. (Wong et al., 2024) noted that the 3 bands of Trojans are similar to spectra of radiolytically processed of C- and N-bearing ices. Such a scenario would be consistent with the production of 3 absorbers from irradiation in the outer solar system prior to Jovian capture, and could explain the absorptions in both red and less-red sufaces. We note an additional possibility, motivated by the discovery of Lysithea’s less-red 3 band shape within the otherwise hydrated Himalia family. Cronstedtite, the Fe-rich serpentine endmember, is the first phyllosilicate to form during aqueous alteration processes (see discussion of alteration progression by McSween et al., 2018). Terrestrial samples show an absorption band centered near to 2.8-2.9 microns (Bates et al., 2020; Gomez Barrientos et al., 2024), in potential agreement to the location of the less-red features. If any irregular satellites contained cronstedtite, that would imply they experienced a minimal but measurable amount of aqueous alteration. These materials are not represented in the meteoritic inventory, but could be analogous to petrographic grade 3 samples. We note that our comparison to Cronstedtite is based on laboratory measurements at ambient conditions (Bates et al., 2020), and contamination from adsorbed water and/or Fe-oxyhydroxides may be present that distort this absorption feature.
8 Ananke: Aqueous Alteration Among Retrograde Satellites?
While Himalia and Elara clearly demonstrate hydrated mineralogies, their bands do not resemble any known meteorites. Similarly, we do not find any clear meteorite analogs for Lysithea and Pasiphae, albeit their 3 micron bands do partially resemble some carbonaceous asteroids observed over the limited wavelength ranges accessible from previous ground-based surveys (Rivkin et al., 2015). These comparisons are severely limited by the small number of laboratory studies of relevant materials that have been prepared at vacuum to remove terrestrially adsorbed water. However, we do find several suitable comparisons relevant to Ananke.
We do not identify a specific spectral analog among the meteorite inventory, but note a broad similarity with a variety of hydrated alteration products found amongst CV, CR, and CM samples.
Fig. 11 shows that Ananke’s band is broadly consistent in shape, depth, and location with some hydrated C2 chondrites. Ananke is intermediate between CR2 Dominion Hills 10085 (Takir et al., 2019) and CM2 Murchison (Yu et al., 2024) in terms of the band center. The broad, rounded feature on Murchison is attributed to the presence of structural OH in phyllosilicates. Ananke also displays a minor absorption at 2.67 and 2.86 microns, along with a weak 3.3-3.4 micron absorption consistent with organic materials. We therefore conclude that Ananke’s 3 micron band is consistent with C2-like phyllosilicate assemblages, and that it differs strongly from the hydrated materials present on Himalia.
By suggesting that Ananke contains phyllosilicates, we provide the first evidence that aqueous alteration amongst the Jovian irregular satellites extends beyond Himalia and its prograde family. Since Ananke is in a retrograde orbital group, its parent body cannot be directly related to Himalia. Combined with Himalia, Ananke’s 3 properties suggests that parent bodies within two of the four main irregular satellite groups experienced aqueous alteration. Both the prograde and retrograde swarms include hydrated asteroids as well as red and less-red Trojan-like bodies.
@@ 9 Conclusions
We report 0.7-5.2 reflectance spectra of eight Jovian irregular satellites collected with JWST’s NIRSpec instrument. We find major variation in composition and presence of aqueous alteration products between each major orbital group. Irregular satellite families, which are presumed to have formed via collisions, contain various Trojan-like and C-type-asteroid-like surfaces throughout both prograde and retrograde populations.
@@
The two largest irregular satellites, Himalia and Elara ( 140 km and 80 km, respectively, both prograde, Himalia group) have bands at 2.7 and 3.05 demonstrating the presence of ammoniated phyllosilicates. Elara’s major band properties are consistent with a simple mean between Himalia and Lysithea. Himalia and Elara also have a minor feature at 4.25 , similar to the feature discovered on the Trojan Eurybates and associated with complexed CO2 (Wong et al., 2024).
The observed spectral variation in the Himalia group, including Elara’s blend of Himalia-like and Lysithea-like material could relate to a stratified mutual parent body. Generally, we suggest that the complex hydration states observed within the Himalia family traces back to a heterogenous parent body, and disfavor exogenic material as the cause of the variability.
@@
We propose that the Himalia parent body formed with materials similar to Ceres-like ammonium-bearing asteroids. Determining the distribution of ammoniated small bodies outside of the main asteroid belt is critical to determine how closely linked Himalia and Ceres may be. We note Himalia is spectrally similar to ground-based observations of the inner satellite Amalthea (Takato et al., 2004), which may represent the primitive materials that accreted near Jupiter prior to its migration. Further compositional characterization of the inner satellites is critical to assess whether Jupiter captured and retained any irregular satellites from zones near to the planet’s formation region.
At least three other distinct types of carbonaceous material are ubiquitous among both prograde and retrograde irregular satellites, with observed matches to both “red” and “less red” Jovian Trojans as well as aqueously altered carbonaceous chondrites. The retrograde satellite Ananke is similar to a variety of hydrated chondrites, suggesting that aqueous alteration is present within the retrograde satellite population and is not limited to the prograde Himalia family.
The Pasiphae family contains objects similar to red and less red Trojan asteroids. Notably, Pasiphae has a 3 absorption feature that is similar to the Himalia family member Lysithea, as well as the Trojan asteroid Eurybates. The origins of this absorber are unclear, and its presence in apparently hydrated (Himalia) and anhydrous (Pasiphae) families complicates a clear association with a single surface type.
The “red” group of satellites (Carme, Sinope, Themisto) display notable differences in their band properties. Carme has a similar band to the less-red Trojan Patroclus, and Carme’s redder color may be related to its deeper 3.4 organics band. The prograde satellite Themisto and the retrograde satellite Sinope have 3.0 and 3.4 bands similar to the red Trojan Leucus. The colors of red objects appear to correlate with the depth of the 3.4 feature. We suggest that the 3 and 3.4 absorbers on “red” irregular satellite surfaces must somehow survive the collisional disruption of their parent bodies, or that such absorbers reform on fresh surfaces faster than disruption can occur.
Ultimately, the variety of surface types and hydration states that we observe suggest that Jupiter captured satellites that include less-red and red Trojans as well as aqueously altered asteroids. This includes the capture of hydrated materials in the Ananke family and the complex materials within the Himalia family itself. The origins of the irregular satellites appear varied and require multiple progenitor populations.
@@ 10 Acknowledgments
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #4028. The authors wish to thank James Bauer, Jessica Sunshine, Kathryn Volk, and Shannon MacKenzie for helpful discussions that improved this work. We also thank Driss Takir for providing the original meteorite spectra used for figure comparisons.

### v2

@@ 7 Trojan-like Surfaces Exist Amongst Fragmented Irregular Satellites
Trojan spectra are reprocessed for methodological consistency and were originally reported and analyzed in Wong et al. (2024).
Previous assessments of the spectrally “red” irregular satellites have relied on color analogies to Jovian Trojan asteroids without specific constraints from absorption features. VNIR measurements (Sharkey et al., 2023; Grav & Holman, 2004) found that irregular satellites outside of the Himalia family map within the color taxonomy of Jovian Trojans. Figure 9 compares the irregular satellites Themisto, Sinope and Carme with the Jovian Trojans Leucus and Patroclus observed by Wong et al. (2024). These JWST observations demonstrate that the reddest irregular satellites indeed share common absorption features at 3.0 and 3.4 with their Trojan counterparts.
Sinope and Themisto both have 3.0 and 3.4 bands that correspond directly to those seen amongst the red Trojans (see Themisto/Sinope and Sinope/Leucus comparisons in Fig. 9). However, Carme, the largest member of its own collisional family, displays a surprising combination of band properties. Carme appears to have a different ratio of organics to hydroxylated minerals than seen in either the redder or less-red Trojan groups, with the organics apparently typical of the redder objects and the hydroxylated minerals typical of the less-red objects. Wong et al. (2024) found that the less-red objects Patroclus and Eurybates display weaker 3.4 organics bands and shorter-wavelength 3.0 bands as compared with the red objects Orus and Leucus. VNIR data discriminates Carme into the “red” taxon and differentiates it from the “less-red” Patroclus (Emery & Brown, 2003; Sharkey et al., 2019). Carme further matches the red Trojan group by having a stronger 3.4 band. But Carme’s 3.0 band clearly differs from the long-wavelength absorption seen on red Trojans, and instead has a shape and central wavelength matched by Patroclus (Fig. 9).
Carme’s band location does not correspond to those seen on any similarly colored objects. Carme’s properties are interesting to consider in the overall context of Trojan color groups. Carme’s spectral properties suggest that spectral slopes of Trojans are controlled by organic abundance and less controlled by the 3-micron absorber. Alternatively, Carme’s band properties themselves could have been altered via collisions, perhaps from implanting exogenous material or from revealing interior materials with differing 3.0 band properties. Determining Carme’s history at this point is difficult, however, without additional context from its other family members’ band properties, as well as a broader understanding of the variation within Trojan band properties.
@@
The presence of Trojan-like 3.0 and 3.4 absorption bands on small Jovian irregular satellites provides new context for the chemical pathways that could have produced them. The irregular satellites’ extensive collisional processing (Bottke et al., 2010) took place after their capture about Jupiter. That the irregular satellites fragmented post-capture requires that their current surfaces were created and evolved within the circumjovian environment. This may differ from similarly sized Trojans, as Trojans larger than likely experienced more collisions in the primordial Kuiper belt than in their current locations (Nesvorný, 2018; Marschall et al., 2022; Bottke et al., 2023). If the 3.0 and 3.4 absorbers were present on irregular satellite parent bodies prior to their capture, then those absorbers must have been present in high enough abundances to survive collisional fragmentation and be incorporated into the surfaces of the present day population. Alternatively, the 3.0 and 3.4 absorbers could be produced in-situ on satellite surfaces via irradiation processing (as discussed by Wong et al., 2024). In-situ surface formation of the 3.0 and 3.4 absorbers would require that they be produced by materials that are stable within the Jovian thermal environment.
8 What causes Less-Red OH Absorptions?
Ananke, Lysithea, and Pasiphae present absorptions with central wavelengths . These absorptions are broadly associated with OH-bearing compounds, but we do not find any specific materials that match them directly. Comparisons to other OH-absorptions on small bodies and meteorites show that the mystery of the absorptions may be related to similar questions on a wide range of solar system surfaces. Specifically, the less-red irregular satellites Ananke, Lysithea, and Pasiphae fill an apparent gap between known Trojans () and any currently identified groups of main belt asteroids (e.g., Rivkin et al., 2022, who grouped objects between those with and those with estimated from ground-based observations).
Figure 10 places these less-red irregular satellite absorptions in comparison with two aqueously altered meteorites (Takir et al., 2019; Yu et al., 2024) and two Trojans (Wong et al., 2024), sorted by their . The comparison meteorites were chosen as they represent some of the longest-wavelength absorptions yet observed on aqueously altered primitive materials, which typically cluster near 2.7-2.8 (Takir et al., 2019). Eurybates and Polymele are chosen as they are representative of the two types of absorptions seen amongst less-red Trojans (Wong et al., 2024; Brown et al., 2025).
The relationships between the spectra in Figure 10 have broad implications for ice-poor surfaces throughout the outer solar system. Within the Jovian system, the prograde Lysithea (Himalia Family) is highly similar to the retrograde satellite Pasiphae (Pasiphae Family), as shown in Fig. 4. The Lysithea/Pasiphae 3 absorber exists within two separate families with dramatically different compositional contexts, suggesting that this material existed on multiple irregular satellite parent bodies. Outside of the Jovian satellite system, Eurybates and Polymele have absorption bands centered similar to the ice-poor Saturnian irregular satellites Albiorix and Siarnaq (Belyakov & Brown, 2025). These Saturnian satellites were themselves noted by Belyakov & Brown (2025) to be similar to the ice-poor Neptune Trojan 2006 RJ103 (Markwardt et al., 2023).
Ananke’s band is centered between the petrologic type-2 chondrites CM2 Murchison (Yu et al., 2024) and CR2 DOM 10085 (Takir et al., 2019). The broad, rounded feature on Murchison is attributed to the presence of structural OH in phyllosilicates. Ananke’s 3 micron band appears similar to this Murchison-like absorption, and we conclude that it is consistent with these known C2 phyllosilicate assemblages. Vilas & Hendrix (2024) report a 0.7 absorption band on one of two collected observations of Ananke caused by phyllosilicates.
Understanding Lysithea’s composition is challenging, as its band center is intermediate between the shorter-wavelength meteorite absorptions and the longer-wavelength Trojan absorptions (see centers indicated in Fig. 10). Previous observations have suggested that Lysithea (and the spectrally similar Pasiphae) contain phyllosilicates (Vilas & Hendrix, 2024), and Lysithea appears related to the otherwise aqueously altered Himalia family. However, Lysithea’s band center is not clearly reconcilable with alteration products from the meteorite record. Some very weakly altered meteorites (petrologic type 3) have minor amounts of phyllosilicates and display absorptions centered (Takir et al., 2013, 2019). However, these meteorite absorptions can be sensitive to major distortions by terrestrial contamination, particularly from Fe-oxyhydroxides. Lysithea’s absorption therefore cannot be entirely understood by analogy to C-chondrite like phyllosilicates. Instead, it is possible that Lysithea contains a mixture of Trojan-like OH compounds and aqueously altered phyllosilicates, which combine to create the shifted band. Such a scenario remains highly speculative, however, as it does not lead to any directly testable spectral parameters at present.
Regardless of the bands’ origins, Lysithea (), Eurybates (), and Polymele () present a possible progression of band centers. It is not clear how closely associated these objects’ compositions could be, but we note that the scale of variation between the two Trojans (Eurybates and Polymele) are approximately the same ( shifts) as the differences between Eurybates and the less-red satellites Lysithea/Pasiphae. Observationally, we do not yet understand if the less-red objects generally present a continuum of 3 bands or if any clustering of spectral parameters exist. Without a broader understanding of the band center variability amongst the Jovian Trojans, it is difficult to further assess the Lysithea/Eurybates band differences. Future characterization of the Eurybates collisional family (Wong et al., 2025) will provide a key opportunity to understand if Lysithea-like absorptions exist amongst the Trojans, or if this band is unique to the Jovian satellites.
Ultimately, the origins of absorptions on less-red Jupiter Trojan and irregular satellite surfaces remain ambiguous. Previously, Wong et al. (2024) noted that the 3 bands of Trojans are generally similar to some spectra of radiolytically processed C- and N-bearing ices, hypothesizing that this evolutionary pathway could fit their detections of Trojan OH absorptions and their nondetections of water ice. The application of the ice irradiation hypothesis to the irregular satellites is more complex, however, as this scenario would require that the OH-bearing weathering product(s) survived collisional resurfacing or were reproduced in the Jovian environment, as discussed in Section 7.
@@ 9 Conclusions
We report 0.7-5.2 reflectance spectra of eight Jovian irregular satellites collected with JWST’s NIRSpec instrument. Despite similar hypotheses for their origins, the irregular satellites are more compositionally diverse than the currently known Jovian Trojans. At least three surface types are present amongst the irregular satellites: Ceres-like surfaces, red-Trojan-like surfaces, and less-red surfaces which are intermediate between Trojans and hydrated asteroids/carbonaceous chondrites.
@@
The two largest irregular satellites have major features similar to those of Ceres-like main-belt asteroids. Himalia and Elara ( 140 km and 80 km, respectively, both in the prograde Himalia family) have bands at 2.7 and 3.05 demonstrating the presence of ammoniated phyllosilicates. Elara’s major band properties are consistent with a simple mean of its family members Himalia and Lysithea. Himalia and Elara also have a minor feature at 4.27 , similar to the feature discovered on the Trojan Eurybates and associated with complexed CO2 (Wong et al., 2024).
@@
We propose that the Himalia parent body formed with materials similar to Ceres-like ammonium-bearing asteroids. It is not clear how closely linked Himalia and Ceres may be without further understanding the distribution of ammoniated materials in small body populations. We note Himalia is spectrally similar to ground-based observations of the inner satellite Amalthea (Takato et al., 2004), which may represent the primitive materials that accreted near Jupiter prior to its migration. Further compositional characterization of the inner satellites is critical to assess whether Jupiter captured and retained any irregular satellites from zones near to the planet’s formation region.
The observed spectral variation in the Himalia group, including Elara’s blend of Himalia-like and Lysithea-like material, could relate to a stratified mutual parent body, similar to the hypothesis of Vilas & Hendrix (2024). We suggest that the complex hydration states observed within the Himalia family are more consistent with a heterogenous parent body and disfavor separate parent body origins for the 2.7 and 3.05 absorbers.
In addition to the ammoniated phyllosilicates in the Himalia family, at least two other distinct types of OH-bearing materials are ubiquitous among both prograde and retrograde irregular satellites. Figure 11 summarizes the band properties of these irregular satellites (excluding Himalia and Elara) in comparison to similar measurements of Jovian Trojans (Wong et al., 2024). The less-red irregular satellites (those with smaller VNIR slopes, plotted with lighter shading in Figure 11) have shorter wavelength 3 bands compared to their counterpart Trojans. The red irregular satellites (higher VNIR slopes and darker shading in Figure 11) also display shorter wavelength absorptions compared to their Trojan equivalents, although their bands are less deep and may therefore be more prone to systematic uncertainties in their band centers. We do not attempt to draw any specific trends in the band parameters shown in Figure 11, but note the wide range of colors, band depths, and band centers that are present amongst these materials.
The “red” group of satellites (Carme, Sinope, Themisto) display one notable outlier compared with “red” Trojans. Carme has a similar band to the less-red Trojan Patroclus, but a deeper 3.4 organics band more similar to those seen amongst the red Trojans as discussed in Section 7. The prograde satellite Themisto and the retrograde satellite Sinope have 3.0 and 3.4 bands similar to the red Trojan Leucus. The colors of red objects appear to correlate with the depth of the 3.4 feature. We do not propose specific compositions for the “red” irregular satellites, but we note that they are the products of intense collisional evolution after their capture about Jupiter (Bottke et al., 2010). Therefore the “red” 3.0 and 3.4 absorbers must either form within the circumjovian environment or represent material that survives major resurfacing events.
The less-red irregular satellites defy simple classification, with bands that are similar to the less-red Trojans but at measurably shorter wavelengths. The retrograde satellite Ananke is similar to some hydrated chondrites, suggesting that aqueous alteration is present within the retrograde satellite population and is not limited to the prograde Himalia family. Pasiphae (Pasiphae Family) has a 3 absorption feature that is similar to Lysithea (Himalia Family). The origins of this absorber are unclear, and its presence amongst two separate families complicates a clear association with a single surface type. We speculate that this group may contain Trojan-like materials that have experienced some weak aqueous alteration to modify their OH absorption bands.
We suggest two possible methods to explain the overall variation of the irregular satellites and their distinctions compared to the Jovian Trojans. The first possibility is that the Himalia and/or Ananke families represent separate compositional lineages that were not captured into the Jovian Trojans. In this case, the capture of the Himalia and Ananke parents would be unrelated to the scattering of the primordial Kuiper belt (the hypothesized origins of the Trojans). Instead, Himalia and Ananke could represent captures from an earlier epoch of Jupiter’s history. This hypothesis can be tested by understanding the compositions of Jupiter’s inner satellites, whose current surfaces may shed further light on the materials that accreted near to Jupiter’s formation region.
Alternatively, for Himalia and Ananke to be related to Trojan compositions, they could represent the unseen cores of Trojan-like objects. A Trojan/Himalia relationship may seem implausible due to the lack of any signs of ammoniation amongst the Jovian Trojans. Similarly, Ananke’s lower wavelength 3 band is unlike any that has been observed on a Trojan. However, the possible presence of Trojan-like material (Lysithea) within the Himalia family suggests that some relationship between Himalia and Trojan-like materials may exist. Connecting either or Himalia to Trojan-like compositions requires that Lysithea-like OH absorptions are directly related to those of the less-red Trojans, despite their different wavelength centers.
Either origin scenario for Himalia and Ananke remains speculative without understanding the nature of variation among any other asteroid collisional families. Future tests for 3 variability among collisionally evolved Trojans (Wong et al., 2025) and main-belt asteroids (Takir et al., 2024) will be useful to explore compositional stratification amongst other populations. Similarly, the origins for Himalia and Ananke could be assessed by understanding if Trojans demonstrate surface- variable 3 bands, particularly surrounding impact craters. The Himalia family demonstrates that surfaces with similar colors can nonetheless have different OH-bearing materials, so such searches should not rely only on color parameters.
@@ 10 Acknowledgments
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #4028. The specific observations analyzed can be accessed via https://doi.org/10.17909/pwcf-c575 (catalog doi: 10.17909/pwcf-c575). The authors wish to thank James Bauer, Jessica Sunshine, Kathryn Volk, and Shannon MacKenzie for helpful discussions that improved this work. We also thank Driss Takir for providing the original meteorite spectra used for figure comparisons.